D is a point on the side BC of an equilateral triangle ABC such that DC is equal to 1/4 of BC. prove that AD square is equal to 13 CD esquare
Answer by mananth(16949)

.
AD^2= AE^2+DE^2
AE= sqrt(3)AC/2
AE^2= 3AC^2/4
AC=BC
AE^2 = 3BC^2/4
But BC = 4DC
AE^2 = 48DC^2/4
AE^2= 12DC^2
DE^2= (CE-CD)^2
But CE = BC/2
But BC= 4DC
SO CE = 2DC
DE^2= (2DC -DC)^2
= DC^2


SO AD^2 = 12CD^2 +DC^2
=13DC^2
AD^2 =13CD^2